# ReDDiT: Rehashing Noise for Discrete Visual Generation

### v2

@@ 3.2 Discrete Diffusion with Rehashing Noise
To exploit visits across all the possible paths, we rewrite the transition kernel defined by Eq. 1 into
@@
We introduce the essential transition matrix to model the process. Let $\mathbf{I}_{(d,m)}$, $\mathbf{M}_{(d,m)}$ and $\bm{\pi}_{(d,m)}$ be matrices in $\mathbb{R}^{(d+m)\times(d+m)}$, defined as
@@
where $\mathbf{I}_{d}$ is the $d\times d$ identity matrix, and $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of ones. The transition matrix $Q_{t|s}\in\mathbb{R}^{(d+m)\times(d+m)}$ corresponding to Eq. 8 is
@@
Rehash Sampling.
To generate a sequence of length $L$, the reverse process starts with $x_{1}\sim\text{U}(\mathcal{M}_{m}^{L})$. Let $\mathbf{a}\odot\mathbf{b}$ denote the Hadamard product between two vectors $\mathbf{a}$ and $\mathbf{b}$, the reverse process is inferred from D3PM deduction [1] as:
[…]
@@
Denote $\alpha_{t|s}^{\rightarrow}=\frac{1-\alpha_{s}}{1-\alpha_{t}}$ to describe the reverse transition rate, and we have
@@
$q(x_{s}^{i}\mid x_{t}^{i})=\begin{cases}1,&\text{if }x_{s}^{i}=x_{t}^{i},\ x_{t}^{i}\notin\mathcal{M}_{m},\\ \alpha_{t|s}^{\rightarrow}/m,&\text{if }x_{s}^{i}\in\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 1-\alpha_{t|s}^{\rightarrow},&\text{if }x_{s}^{i}\notin\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 0,&\text{otherwise.}\\ \end{cases}$ (17)
@@
$p_{\theta}^{i}(x_{s}^{i}|x_{t})=\begin{cases}1,&\text{if }x_{s}^{i}=x_{t}^{i},\ x_{t}^{i}\notin\mathcal{M}_{m},\\ \alpha_{t|s}^{\rightarrow}/m,&\text{if }x_{s}^{i}\in\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ (1-\alpha_{t|s}^{\rightarrow})p_{\theta}^{i}(x_{t}),&\text{if }x_{s}^{i}\notin\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 0,&\text{otherwise.}\\ \end{cases}$ (18)
The subsequent latents $x_{t}$ are generated by discretizing the reverse timeline $T$ into $K$ steps. We denote this schedule as $T^{1:K+1}$ such that $T^{1}=1$ and $T^{K+1}=\varepsilon$, with $\varepsilon$ being an arbitrarily small positive constant. Inspired by MD4 [36], we use non-uniform time intervals for efficient sampling.
Algorithm 1 MVTM Sampling
1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$,
2:Settings: codebook entries $\mathcal{V}_{(d,1)}$, number of steps $K$, $G(t)$, Gumbel noise $\mathcal{G}$
3:Initialize: $x_{1}\leftarrow\mathcal{M}_{1}^{L}$, $t\leftarrow 1$.
4:for $k=1$ to $K$ do
5:     $t\leftarrow\frac{K-k+1}{K},s\leftarrow\frac{K-k}{K}$
6:     $\text{logits}\leftarrow f_{\theta}(x_{t},c)$
7:     $p_{\text{score}}\leftarrow\text{logits}+G(t)\cdot\mathcal{G}$
8:     $x_{\text{pred}}\leftarrow\text{argmax}(p_{\text{score}})$triangleright Predict-all
9:     $x_{s}\leftarrow\text{where}(x_{t}=[\text{m}],x_{\text{pred}},x_{t})$
10:     $p_{\text{conf}}\leftarrow p_{\text{score}}+G(t)\cdot\mathcal{G}$
11:     $m_{\text{re}}\leftarrow\text{argsort}(p_{\text{conf}})[1:L\cdot(1-\alpha_{s})]$
12:     $x_{s}\leftarrow\text{where}(m_{\text{re}},[\text{m}],x_{s})$triangleright Re-mask
13:end for
14:Return: fully unmasked sequence $x_{0}$
Algorithm 2 Rehash Sampling (Ours)
1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$.
2:Settings: codebook entries $\mathcal{V}_{(d,m)}$, number of steps $K$.
3:Initialize: $x_{1}\leftarrow\text{U}(\mathcal{M}_{m}^{L})$, $t\leftarrow 1$, $T^{1:K}$.
4:for $k=1$ to $K$ do
5:     $t\leftarrow T^{k},s\leftarrow T^{k+1}$
6:    # the rehash operation:
7:     $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$
8:     $\text{logits}\leftarrow f_{\theta}(x_{t},c)$
9:     $p_{0|t}\leftarrow\text{Softmax}(\text{logits})$
10:     $q_{s|t}\leftarrow\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\cdot p_{0|t}+\delta_{m[0]}\cdot\frac{1-\alpha_{s}}{1-\alpha_{t}}$
11:     $x_{\text{pred}}\leftarrow\text{Multnm.}(q_{s|t})$triangleright w/ masks
12:     $x_{s}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},x_{\text{pred}},x_{t})$
13:end for
14:Return: fully unmasked sequence $x_{0}$
In practice, we apply torch.multinomial (Multnm. in step 10) for low-discrepancy categorical sampling [18] with the merged probability of absorbing states  [^1] Instead of dividing $\alpha_{t|s}^{\rightarrow}$ and assigning these probabilities to each mask vocabulary. We merge the probabilities at step 9 in Alg. 2 to keep an overall noise sampling probability, as small values might be truncated.. The random nature of absorbing states inspires a rehash operation: we shuffle these tokens at the beginning of each step by $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$.
3.3 Discussion
Comparison with MVTM.
Masked visual token models (MVTMs) borrow the objective
from masked language models [7] and predict on masked tokens with a maximum likelihood. Besides the reformulated corruption (Eq. 8) and reverse process (Eq. 17), ReDDiT differs in the following aspects: (i) the training objective (Eq. 4), which is derived from the principled DDM which provides better theoretical and empirical results with rehashing noise design. (ii) it can easily sample with arbitrarily discretized timeline, while MVTM couples training and inference, restricting its flexibility; (iii) the rehash sampler (Alg. 2) includes the absorbing states in categorical sampling with lower discrepancy, different from predict-remask sampler with time variant intensity $G(t)$ over Gumbel noise $\mathcal{G}$ (Alg. 1)  [^2] The logits corresponding to previously restored tokens’ indices are manually set to infinity for both methods, so that they will not be noised again in the following steps. This leads to an implementation of any-order auto-regressive model [26] if DDM’s decoded tokens per step is limited to $1$.. Gumbel-max suffers from numerical inaccuracy [49] and we detect that it worsens on large vocabulary (Fig. 3 with our reproduced results), which limits the potential of MVTM.
Relationship to DFM.
Discrete flow matching (DFM) [11] applies flow matching techniques to discrete structures and introduces a transition process based on masked tokens. Its training objective was initially designed as the masked token loss (19), and evolved to a time-weighted cross-entropy loss [34] for generalized diffusion paths, which is similar to ours. The similarity enables a direct comparison between the DFM sampler and our rehash sampler using the same trained model weights. We noticed that while the DFM sampler offers a refinement mechanism via token-wise updates, it generally requires more steps to reach optimal results. However, since the gradual decoding method is shared, we can integrate certain DFM steps into our sampling procedure for refinement. We leave this topic for an updated version of our essay.
Figure 3: Comparison of methods under limited samples. Left: Gumbel-max is theoretically equivalent to our method, yet it struggles to reflect the true distribution under limited sample passes. The multinomial approach captures the distribution more accurately. Right: our model achieves lower gFID across different sampling steps without tuning Gumbel-max, indicating more efficient and faithful sampling. (Best viewed in color)
4 Experiment
4.1 Implementation
Datasets.

### v3

@@ 2.2 Discrete Diffusion with Rehashing Noise
To exploit visits across all the possible paths, we rewrite the transition kernel defined by Eq. 1 as
$q(x_{t}^{i}|x_{s}^{i})=\begin{cases}1-\alpha_{t|s},&\text{if }x_{t}^{i}\in\mathcal{M}_{m},\ x_{s}^{i}\notin\mathcal{M}_{m}\\ \alpha_{t|s},&\text{if }x_{t}^{i}=x_{s}^{i},\ x_{s}^{i}\notin\mathcal{M}_{m}\\ 1/m,&\text{if }x_{t}^{i}\in\mathcal{M}_{m},\ x_{s}^{i}\in\mathcal{M}_{m}\\ 0,&\text{otherwise}.\end{cases}$ (8)
With above definitions, we reformulate the diffusion process of $x$ as a transition from $\mathcal{E}_{d}$ to $\mathcal{M}_{m}$. We train the model by feeding it with corrupted data, of which the distribution is inferred as $x_{t}\sim\text{Cat}(x_{t};\alpha_{t}x_{0}+(1-\alpha_{t})\text{U}(\mathcal{M}_{m}^{L}))$, where $\text{U}(\mathcal{M}_{m}^{L})$ is the uniform distribution upon $\mathcal{M}_{m}^{L}$.
Rehash Sampling.
To generate a sequence of length $L$, the reverse process starts with $x_{1}\sim\text{U}(\mathcal{M}_{m}^{L})$. The subsequent latents $x_{t}$ are generated by discretizing the reverse timeline $T$ to $K$ steps. We denote this schedule as $T^{1:K+1}$ such that $T^{1}=1$ and $T^{K+1}=\varepsilon$, with $\varepsilon$ being an arbitrarily small positive constant. The reverse process is deduced from the formulation, as
[…]
Comparing with MVTM sampler in Alg. 1, our rehash sampler is shown in Alg. 2. Similar to MDLM Sahoo et al. (2024), we apply torch.multinomial (Multnm. in step 10) for low-discrepancy [^1] Instead of dividing $\alpha_{s|t}$ and assigning these probabilities to each mask vocabulary. We merge the probabilities at step 9 to keep an overall noise sampling probability, as small values might be truncated. categorical sampling.
Algorithm 1 MVTM Sampling
1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$,
2:Settings: number of steps $K$, $G(t)$, $\mathcal{G}$
3:Initialize: $x_{1}\leftarrow\mathcal{M}_{1}^{L}$, $t\leftarrow 1$.
4:for $k=1$ to $K$ do
5:  $t\leftarrow\frac{K-k+1}{K},s\leftarrow\frac{K-k}{K}$
6:  $p_{\text{score}}\leftarrow f_{\theta}(x_{t},c)+G(t)\cdot\mathcal{G}$
7:  $x_{\text{pred}}\leftarrow\text{argmax}(p_{\text{score}})$ $\triangleright$ Predict-all
8:  $x_{s}\leftarrow\text{where}(x_{t}=[\text{m}],x_{\text{pred}},x_{t})$
9:  $p_{\text{conf}}\leftarrow p_{\text{score}}+G(t)\cdot\mathcal{G}$
10:  $m_{\text{re}}\leftarrow\text{argsort}(p_{\text{conf}})[1:L\cdot(1-\alpha_{s})]$
11:  $x_{s}\leftarrow\text{where}(m_{\text{re}},[\text{m}],x_{s})$ $\triangleright$ Re-mask
12:end for
13:Return: fully unmasked sequence $x_{0}$
Algorithm 2 Rehash Sampling
1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$.
2:Settings: number of steps $K$.
3:Initialize: $x_{1}\sim\text{U}(\mathcal{M}_{m}^{L})$, $t\leftarrow 1$, $T^{1:K}$.
4:for $k=1$ to $K$ do
5:  $t\leftarrow T^{k},s\leftarrow T^{k+1}$
6:  $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$
7:  $p\leftarrow\pi_{\theta}(x_{t},c)$
8:  $q_{s|t}\leftarrow\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\cdot p+\delta\cdot\frac{1-\alpha_{s}}{1-\alpha_{t}}$
9:  $x_{\text{pred}}\leftarrow\text{Multnm.}(q_{s|t})$ $\triangleright$ w/ masks
10:  $x_{s}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},x_{\text{pred}},x_{t})$
11:end for
12:Return: fully unmasked sequence $x_{0}$
The random nature of absorbing states inspires a rehash operation: we shuffle these tokens at the beginning of each step by $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$. Proof to Eq.9 is included in Appendix. C.
2.3 Discussion
Figure 3: Sampler comparison. Left: Gumbel-max is theoretically equivalent to our method, yet it struggles to reflect the true distribution under limited sample passes. The multinomial approach captures the distribution more accurately. Right: our model achieves lower gFID across different sampling steps without tuning Gumbel-max, indicating more efficient and faithful sampling. a, b, c refer to three uniformly sampled $G(t)$ set for MVTM sampling. See supplementary for experimental codes. (This figure is best viewed in color)
Comparison with MVTM.
Masked visual token models (MVTMs) borrow the objective
from masked language models Devlin et al. (2019) and predict on masked tokens with a maximum likelihood. Besides the reformulated corruption (Eq. 8) and reverse process (Eq. 9), ReDDiT differs in the following aspects: (i) the training objective (Eq. 4), which is derived from DDM, providing better theoretical and empirical results. (ii) it can easily sample with a arbitrarily discretized timeline, while MVTM couples training and inference, restricting its sampling flexibility; (iii) the rehash sampler (Alg. 2) includes absorbing states in categorical sampling with lower discrepancy, different from MVTM’s predict-remask sampler with time variant intensity $G(t)$ over Gumbel noise $\mathcal{G}$ (Alg. 1)  [^2] The logits corresponding to previously restored tokens’ indices are manually set to infinity for both methods, so that they will not be noised again in the following steps. This leads to an implementation of any-order auto-regressive model Ou et al. (2024) if DDM’s decoded tokens per step is limited to $1$.. Gumbel-max suffers from numerical inaccuracy Zheng et al. (2024) and we noitice that it becomes worse on large vocabulary (Fig. 13 with our reproduced results), which limits MVTM’s potential.
Relationship to DFM.
Discrete flow matching (DFM) Gat et al. (2024) introduces a transition process based on masked tokens. Its training objective was initially designed as the masked token loss ( 10), and evolved to a time-weighted cross-entropy loss Shaul et al. (2024) for generalized diffusion paths, which is similar to ours. The similarity enables a direct comparison between the DFM sampler and our rehash sampler using the same trained model weights. We notice that it generally requires more steps to reach optimal results, as the DFM sampler offers a refinement mechanism via token-wise updates. Since the gradual decoding method is shared, we can integrate certain DFM steps into our sampling procedure for refinement. This leads to $\sim$ 0.1 gFID improvement on ImageNet-1K. Refer to Appendix D for details.
3 Experiment
3.1 Implementation
Datasets.
[…]
Pre-processing.
Following the setting in LlamaGen Sun et al. (2024), we apply the ten-crop augmentation on images, and use pre-trained tokenizers to convert them to discrete tokens. We pick IBQ-f16 Shi et al. (2025) tokenizer as default for its scalable and promising performance in generation tasks, which uses a $16\times 16$ downsampling ratio and converts a $256\times 256$ image into 256 discrete tokens. The tokenizer has a codebook with 16384 entries. The LlamaGen-f16 (used in Tab. 2) and LlamaGen-f8 tokenizer Sun et al. (2024) (used in Tab. 1) are also used for comparison with recent discrete generation methods. All tokenizers are used out-of-the-box without modification.
Representation Alignment.
Recent study Yu et al. (2025) has shown that the alignment of intermediate representations between diffusion transformers and vision encoders accelerates training convergence of diffusion models. Accordingly, the alignment is designed as a regularization term with $\lambda=0.5$. We extract diffusion transformer’s 8-th layer intermediate feature $\textbf{h}^{[n]}(x_{t})$ and align it with the original image’s dinov2-b Oquab et al. (2023) encoded features $f_{\text{enc}}(x_{0}^{\text{ori}})$. The intermediate features are projected by a small trainable MLP $h_{\varphi}$. The $\text{sim}(\cdot,\cdot)$ computes the mean of element-wise cosine similarity between embeddings, as
This alignment was proposed for continuous diffusion models, and we firstly validate that it’s also suitable for training discrete models. However, from our observation, as a training acceleration technique, RepA does not provide relative performance gain if training sufficiently (like for 1M steps as most diffusion models do) for discrete latents. We only use RepA to improve training efficiency and probe the inner dynamics through training as in Fig. 4.
Training and Evaluation.
The proposed model is based on DiT Peebles & Xie (2023) architecture, with reference to its discrete prediction version Sahoo et al. (2024). 2D-RoPE Su et al. (2024) and min-SNR Zhang & Sennrich (2019) are applied for training efficiency. The model is optimized using the AdamW optimizer with a cosine decay. Training is conducted for 500k iterations on 8 NVIDIA H100 GPUs with a global batch size 1024. Class-conditional training is enabled using class embeddings and a drop-rate of 0.1 for generation with classifier-free guidance. Details are provided in Appendix E.
3.2 Performance and Comparison
We compare the proposed ReDDiT model with other generative models on the ImageNet-1K $256\times 256$ in Tab. 1. The IBQ tokenizer is used for the default L and XL models. We also utilize LlamaGen-f8 with 128 noise capacity to evaluate its high-resolution potentials (noted as ReDDiT-XLf8). We use a linear increasing guidance following the common practice of  Gao et al. (2023).
Table 1: Performance comparison on class-conditional ImageNet $\times256$. Look-up free quantizers are beyond the scope of this paper. ft.(in gray) indicates that the decoder is fine-tuned for quantized latents. Wall-clock inference time relative to ReDDiT-XL is reported.
Type Model Tokenizer Generator
[…]
Diff. LDM-4 Rombach et al. (2022a) 4096×3 - 3.60 247.7 400M 250
DiT-XL/2 Peebles & Xie (2023) 1024×4 - 2.27 278.2 675M 250 18
SiT-XL Ma et al. (2024) 1024×4 - 2.42 238.5 675M 30 2
SiT-XL w/ Solver Wang et al. (2025) 1024×4 - 2.24 244.1 730M 15 1.2
AR Taming-VQGAN Esser et al. (2021) 256 1024 15.78 74.3 1.4B 256 8
RQ-Transformer Huang et al. (2023) 256 16384 7.55 134.0 3.8B 64 8.5
ViT-VQGAN Yu et al. (2022) 1024 8192 4.17 175.1 1.7B 1024 >10
LlamaGen-3B Sun et al. (2024) 576 16384 2.18 263.3 3.1B 576 20
RandAR-XXL Pang et al. (2024) 512 16384 2.15 322.0 1.4B 88 4
VAR-d30 Tian et al. (2024) 680 4096 1.97 334.7 2.0B 10 0.5
MVTM MaskGIT Chang et al. (2022) 256 1024 6.18 182.1 227M 8 0.2
MaskGIL-XXL Xin et al. (2025) 256 16384 3.71 303.4 1.4B 8 0.8
TiTok-S-128ft. Yu et al. (2024) 128 4096 1.97 281.8 287M 64 1.6
DDM ITM Hu & Ommer (2024) 1024 16384 5.30 183.0 546M 100 3
ReDDiT-L (ours) 256 16384 2.13 294.7 346M 20 0.5
ReDDiT-XL (ours) 256 16384 1.74 313.6 675M 32 1
ReDDiT-XLf8 (ours) 1024 16384 1.61 318.5 675M 64 2
Model VQ Tokenizer Info. Generator
[…]
LlamaGen-LAR Sun et al. (2024) LlamaGen-f16 Sun et al. (2024) 2.19 8 343M 3.80
RandAR-LAR Pang et al. (2024) 343M 2.55
OursDDM(ReDDiT-L) 346M 2.41
IBQ-BAR Shi et al. (2025) IBQ-tokenizer Shi et al. (2025) 1.37 256 343M 2.88
OursDDM(ReDDiT-L) 346M 2.13
Generation Quality.
As shown in Tab. 1, ReDDiT achieves the best performance among the compared discrete models. It outperforms the baseline (MaskGIT Chang et al. (2022)) with significant margins (gFID: 2.13 vs 6.18 and IS: 294.7 vs. 182.1). It also outperforms the recent DDM method Hu & Ommer (2024) and TiTok-S-128 Yu et al. (2024), which is extensively fine-tuned on quantized latents. Compared with continuous diffusion models, ReDDiT exhibits on-par efficiency and performance, showing great potential for discrete generation. Note that the performance is achieved with a codebook size of 16384, validating ReDDiT’s effectiveness for large-vocabulary codebooks.
Efficiency.
ReDDiT is born with the high-efficiency advantage of discrete diffusion models, comparing with AR models. As shown in Tab. 1, the inference time of ReDDiT is slightly longer than MaskGIT, while the performance is overwhelming. Without acceleration techniques, ReDDiT achieves a competitive performance which AR and traditional diffusion models use more than 250 steps to achieve. Notably, when armed with recent efforts that tailored KV-Cache Liu et al. (2025) for discrete diffusion models, ReDDiT’s inference can be further boosted (not included in the main paper for fair comparison). See Appendix F for details.
Besides the major comparison, we also conduct an experiment that utilizes the identical tokenizer in previous AR models and validate our method’s effectiveness. As can be seen in Tab. 2, ReDDiT outperforms AR methods in generation tasks across different tokenizers. Note that this comparison is to demonstrate diffusion model’s potential on discretized latents, and current representation alignment methods are inapplicable to AR models due to their unidirectional attention design.
3.3 Determining Noise Capacity
Figure 4: Comparison of noise capacities. We re-implemented training with MVTM and ReDDiT design with the same training recipe (LlamaGen-f16 as visual tokenizer, with codebook size $16384$). The generation quality and representation alignment trends are visualized.
The reformulated discrete diffusion dynamics defines transitioning from $\mathcal{E}_{d}$ to $\mathcal{M}_{m}$. Under this setting, it is necessary to empirically determine the optimal value of $m$ for a fixed tokenizer with vocabulary size $d$, as the latent representations learned by VAEs are variant. We keep the training setup fixed and conduct experiments w.r.t. the noise capacity $m$. We also visualize $\mathcal{L}_{\text{RepA}}$, which captures the degree of representation alignment Yu et al. (2025) within the transformer.
3.4 Ablation Study
[…]
Sampling Timeline.
Recovering complete information from noise remains critical to diffusion-based models Lu et al. (2022); Wu et al. (2024). Recent work shows MVTM’s non-linear scheduler for training is less critical when using high-capacity tokenizers. Evidence of time-invariance in DDMs Sahoo et al. (2024); Shi et al. (2024) further supports decoupling training from sampling. In our experiments, a linear scheduler with constant signal-to-noise ratio decay, yields optimal training dynamics. Among the timeline discretization tested, Fig. 5, the cosine schedule is employed for our ReDDiT model for best performance in Tab. 3.
General Design.
(a) (a) General model design
Train Config Sample Config gFID Prec. Rec.
MVTM + RepA loss MVTM sampler 6.83 0.75 0.39
Switch to objective (11) MVTM sampler 6.23 0.77 0.41
   same as above Rehash sampler 5.75 0.78 0.45
+ 2D-RoPE + min-SNR Rehash sampler 5.51 0.79 0.45
   same as above + DFM refine 5.40 0.81 0.52
(b) (b) Sampling timeline
Steps Timeline gFID
20 linear 7.18
32 linear 6.43
20 arccos 5.04
20 square 7.39
20 cosine 4.91
We ablate the general choices of ReDDiT, which starts with a re-trained MVTM baseline methods (with LlamaGen-f16 and RepA for faster convergence as default) in Tab. 3. The applied techniques like 2D-RoPE are also ablated with re-training. As shown, through the revised objective and our proposed sampler, ReDDiT alone improves FID by $\sim 1.0$ compared to the baseline model. When combined with modern modification on transformers, it can further improve the performance, showing its complementaryness with main-stream efforts.
3.5 Qualitative Result
Class-conditional Generation.
Figure 6 presents representative class-conditional samples generated by the proposed ReDDiT model. The outputs across diverse image classes consistently exhibit high fidelity and diversity. Additional qualitative comparisons and more sample visualizations are provided in Appendix G.
Image Editing.
We further demonstrate ReDDiT’s editing capability in Figure 6, highlighting its bi-directional perceptual competence. Following MaskGIT Chang et al. (2022), we replace a region of the input image with noise tokens and employ the same generation pipeline to inpaint the missing content, conditioned on a class label $c$. Thanks to the rehashing noise mechanism, ReDDiT is able to produce diverse and semantically coherent completions without adjusting temperature or other sampling parameters.
4 Conclusion
We proposed ReDDiT, a discrete visual generative model built upon a discrete diffusion architecture with novel noise designs and efficient sampling strategies. Our key contribution lies in the integration of rehashing noise with samplers, which together ensure both diversity and low discrepancy throughout the generative process. By introducing rehashing noise, ReDDiT enriches the potential paths that latent variables can traverse during training, regularize training dynamics and enhances model’s representational capacity. Extensive experiments demonstrate that discrete generative models can achieve performance on par with their continuous counterparts while offering top-tier efficiency. This study paves a promising way for discrete generative modeling and offers fresh insights toward unifying visual and language generation—a path we leave for future exploration.
Appendix A Use of Large Language Models
In the process of drafting this paper, Large Language Models (LLMs) were solely utilized for writing polish (e.g., optimizing sentence structure, enhancing expression fluency). No LLM was involved in core academic work such as conceptualization, literature review, data analysis, argument construction, or conclusion formulation of this study.
As the human authors of this paper, we bear full and sole responsibility for the paper’s content, including the accuracy of research data, validity of academic arguments, integrity of research methods, and compliance with academic ethics.
Appendix B Related Work
Diffusion Models.
Diffusion models Ho et al. (2020); Song et al. (2020) have emerged as a powerful class of generative methods that learn data distributions by reversing a gradual noising process over time. These models are primarily designed for continuous domains such as images Dhariwal & Nichol (2021); Gao et al. (2023); Peebles & Xie (2023), defining a forward process that transforms data $x_{0}$ into noise $x_{1}$: $x_{t}\sim\mathcal{N}(\sqrt{\alpha_{t}}x_{0};(1-\alpha_{t})\mathbf{I})$ where $\alpha_{t}$ controls the noise schedule. The generative (reverse) process learns a denoising model $p_{\theta}(x_{s}\mid x_{t})$, often parameterized via a neural network $\theta$ to predict either noise or clean data.
Discrete Diffusion Models.
Discrete diffusion has been previously governed by masked visual token models (MVTMs) Chang et al. (2022; 2023); Gu et al. (2022); Yu et al. (2023; 2024). This model leverages a BERT-style [mask] token to corrupt the tokenized image sequence and trained the network with a simple cross-entropy loss on masked tokens, resulting in a score-based prediction. It generates tokens in a non-autoregressive fashion, by remasking the tokens with least scores at each inference as depicted in Alg. 1.
Recent studies unlocked the principled discrete diffusion model (DDM) Sahoo et al. (2024); Shi et al. (2024) and discrete flow-matching (DFM) Gat et al. (2024); Shaul et al. (2024), which adapt the Markov chain theory, enabling generation over text Ou et al. (2024); Nie et al. (2025), moleculesShaul et al. (2024), and other discrete representations Austin et al. (2021); Nisonoff et al. (2024). Unlike MVTMs, the principled DDM and DFM mostly derive a time-weighted cross-entropy loss to supervise the training procedure and apply a gradual unmasking method based on probabilities.
Appendix C Discrete Diffusion with Rehashing Noise
Complete Definition and Deduction.
We provide a full theoretical discussion on the corrupted distribution and reverse process defined in the main paper. The extended discussions with corresponding proofs are marked with teal.
Given $d$ categories, let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be its one-hot vector where the $i$-th value is $1$. We denote $\mathcal{E}=\{\mathbf{e}_{i}\in\mathbb{R}^{d}\mid i=1,\ldots,d\}$ as the basis of a categorical distribution, and a basis for absorbing states with capacity $m$: $\mathcal{M}=\{\mathbf{m}_{j}\in\mathbb{R}^{m}\mid j=1,\ldots,m\}$. The sum of $\mathcal{E}$ and $\mathcal{M}$ can be denoted as
To exploit visits across all the possible paths, for $0\leq s<t\leq 1$, we write the transition kernel as [^4] To maintain simplicity, we use $\alpha_{t|s}^{\leftarrow}=\frac{\alpha_{t}}{\alpha_{s}}$ and $\alpha_{t|s}^{\rightarrow}=\frac{1-\alpha_{s}}{1-\alpha_{t}}$ to denote transition rate for the corruption and reverse process, respectively.
@@ Appendix C Discrete Diffusion with Rehashing Noise
Proof of the Corrupted Distribution.
The presentation in the main paper simplifies the theory without specifying the transition matrix $Q_{t}$ due to page limitation. We make a detailed version with important yet basic matrix calculation in this section.
Let $\mathbf{I}_{(d,m)}$, $\mathbf{M}_{(d,m)}$ and $\bm{\pi}_{(d,m)}$ be matrices in $\mathbb{R}^{(d+m)\times(d+m)}$, defined as
@@
where $\mathbf{I}_{d}$ is the $d\times d$ identity matrix, and $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of ones.
The transition matrix $Q_{t|s}\in\mathbb{R}^{(d+m)\times(d+m)}$ is defined as:
@@
Proof of the Reverse Process.
To generate a sequence of length $L$, the reverse process starts with $x_{1}\sim\text{U}(\mathcal{M}_{m}^{L})$. Let $\mathbf{a}\odot\mathbf{b}$ denote the Hadamard product between two vectors $\mathbf{a}$ and $\mathbf{b}$, the reverse process is inferred as:
$q(x_{s}\mid x_{t})$ $=\dfrac{Q_{t|s}x_{t}\odot Q_{s|0}^{\top}x_{0}}{x_{t}^{\top}Q_{t|0}^{\top}x_{0}}\ \ \ \ \text{(D3PM deduction)}$
@@
Notice that $\alpha_{t|s}^{\rightarrow}=\frac{1-\alpha_{s}}{1-\alpha_{t}}$, and we have
@@
[…]